# Deep Linear Network Training Dynamics from Random Initialization: Data, Width, Depth, and Hyperparameter Transfer

### v2

@@ 2 Deep Linear Network with Random Data
$f(\bm{x})=\frac{1}{N\gamma_{0}}\left(\bm{w}^{L}\right)^{\top}\prod_{\ell=1}^{L-1}\left(\frac{1}{\sqrt{N}}\bm{W}^{\ell}\right)\left(\frac{1}{\sqrt{D}}\bm{W}^{0}\right)\bm{x}.$ (1)
The first matrix has dimensions $\bm{W}^{0}\in\mathbb{R}^{N\times D}$, while all intermediate weight matrices are all of size $\bm{W}^{\ell}\in\mathbb{R}^{N\times N}$. Lastly, the output $\bm{w}^{L}\in\mathbb{R}^{N}$ maps to a single output. We can provide extensions to unmatched hidden widths, structured data, and multiple (finite) output classes in Appendix LABEL:app:diff_rel_widths,_app:structured_data,_app:multiple_outputs, but we begin our analysis with this simplest possible case. We start by characterizing the dynamics for a random training dataset $\mathcal{D}=\{\bm{x}_{\mu},y_{\mu}\}_{\mu=1}^{P}$ of size $P$, and random initial weights
@@ 2.1 Dynamics of (S)GD in a Joint Asymptotic Limit
$\bm{v}(t)=\bm{w}^{\star}-\frac{\sqrt{D}}{N\gamma_{0}}\bm{W}^{1}(t)^{\top}\bm{g}^{1}(t).$ (9)
From the correlation functions, we can directly obtain the test loss $\mathcal{L}(t)$ and train loss ^$\hat{\mathcal{L}}(t)$ as
@@
Figure 2: The theory can capture the failure of hyperparameter transfer in NTK parameterization and the success of hyperparameter transfer in $\mu$P networks. (a) The optimal learning rate increases with model width $\nu=N/D$ in NTK parameterization. This is due to the fact that the effective feature learning rate $\gamma_{0}$ is decreasing as $1/\sqrt{\nu}$. (b) The optimal learning rate is approximately preserved across $\nu=N/D$ in $\mu$P networks where $\gamma_{0}$ is constant, enabling learning rate transfer.
@@
(a) Test Loss Varying Data $\gamma_{0}=2$
(b) Emergence of Overfitting from Data Repetition
(c) Online SGD Dynamics $\nu=0.25$
(d) SGD varying $\nu$ at Small $\alpha_{B}=0.05$
Figure 3: Train and test losses of a depth $L=4$ linear network versus dataset size $P=\alpha D$ compared to online SGD with batch size $B=\alpha_{B}D$. (a) Networks can fit the training data faster for smaller datasets $\alpha$. (b) The test error can be non-monotonic as a function of data $\alpha$ if training is performed too long. (c) The dynamics of online SGD with $\nu=0.5$ across batchsizes $B=\alpha_{B}D$. Smaller batchsizes slow down training through accumulated SGD variance, but there is not a bottleneck error as $t\to\infty$ provided the dynamics are stable. (d) Increasing the width $\nu$ eliminates the initial overfitting transient.
@@ 3 Residual Networks and Depth Limits
$f(\bm{x})=\frac{1}{N\gamma_{0}}\left(\bm{w}^{L}\right)^{\top}\prod_{\ell=1}^{L-1}\left(\bm{I}+\frac{\beta}{\sqrt{N}}\bm{W}^{\ell}\right)\left(\frac{1}{\sqrt{D}}\bm{W}^{0}\right)\bm{x}$ (16)
@@ 4 Structured Data & Power Law Scalings
which agrees with the exponents derived by Bordelon et al. (2024b) in the one-hidden layer $L=1$ case. However, our theory shows that this improved exponent for the hard task regime due to non-lazy dynamics is in fact the same for deeper networks as well. This indicates that the structure of the task could induce some universal behavior in linear networks of different architectures.
@@ Discussion
The current theory, while a signficant improvement in complexity compared to prior works which do not average over the random data (such as Bordelon & Pehlevan (2022); Bordelon et al. (2024c)), becomes costly to compute for very deep networks trained for a large number of iterations $T$ and very deep networks $L\gg 1$. We summarize the memory and compute requirements to compute our theory in Table 1. Because of the $\sim L^{2}$ number of response functions required to compute our residual network DMFT, the costs are higher than for a standard non-residual network (which is linear in $L$). Our DMFT solver is more efficient than full batch training in a finite width $N$ linear ResNet provided that $LT^{2}\ll N^{2}P$. In addition, the function approximation power of this model is quite limited as we are only considering deep linear networks in the present study. Some future directions that could be analyzed within this framework include characterizing the impact of learning rate schedules and other optimizers on the dynamics, studying the effect of normalization layers, and characterizing how nonlinearity changes this theoretical picture.
@@ Acknowledgements
We thank Boris Hanin, Mufan (Bill) Li, Lorenzo Noci, Jacob Zavatone-Veth, Hugo Cui, Francesca Mignacco, Stefano Sarao Manelli, and Clarissa Lauditi for useful and engaging conversations.
@@ A.1.1 Isolating Dependence on The Disorder
$f(\bm{x})=\frac{1}{N\gamma_{0}}(\bm{w}^{L})^{\top}\prod_{\ell=1}^{L-1}\left(\frac{1}{\sqrt{N}}\bm{W}^{\ell}\right)\left(\frac{1}{\sqrt{D}}\bm{W}^{0}\right)\bm{x}.$ (22)
@@
$\bm{w}^{L}(t+1)=\bm{w}^{L}(t)-N\eta\gamma_{0}^{2}\nabla_{\bm{w}^{L}}\mathcal{L}(t)\ ,\ \bm{W}^{\ell}(t+1)=\bm{W}^{\ell}(t)-N\eta\gamma_{0}^{2}\nabla_{\bm{W}^{\ell}}\mathcal{L}(t)\ ,\ \ell\in\{0,...,L-1\}.$ (24)
@@ A.2 Final DMFT Equations
$v(t)=w^{\star}-r^{0}(t)-\sum_{t^{\prime}<t}\left[\frac{1}{\gamma_{0}}R^{1}_{gu}(t,t^{\prime})h^{0}(t^{\prime})+\eta\gamma_{0}C_{g}^{1}(t,t^{\prime})\right]h^{0}(t^{\prime})\ ,\ r^{0}(t)\sim\mathcal{GP}\left(0,\frac{1}{\nu\gamma_{0}^{2}}G^{1}\right)$
@@
$h^{1}(t)=u^{1}(t)+\sum_{t^{\prime}<t}\left[\frac{1}{\nu\gamma_{0}}R_{hr}^{0}(t,t^{\prime})g^{1}(t^{\prime})+\eta\gamma_{0}C_{h}^{0}(t,t^{\prime})\right]g^{1}(t^{\prime})\ ,\ u^{1}(t)\sim\mathcal{GP}(0,C_{h}^{0})$
@@ A.3 Centering: Subtracting the Initial Predictor for Lazy Limit Stability
$v(t)=w_{\star}-r^{0}(t)+r^{0}(0)-\frac{1}{\gamma_{0}}\sum_{t^{\prime}<t}R^{1}_{gu}(t,t^{\prime})h^{0}(t^{\prime})-\eta\sum_{t^{\prime}<t}G^{1}(t,t^{\prime})h^{0}(t^{\prime})$
@@ Appendix B Comparing Parameterizations
In this section, we briefly comment on how the maximal stable learning rate scales with $\nu$ in NTK and in mean field/$\mu$P parameterization. NTK parameterization is defined as scaling the network so that $\gamma_{0}=\Theta\left(\frac{1}{\sqrt{\nu}}\right)$. However, a consequence of this choice is that the change in the correlation functions $C_{g}(t,t^{\prime})$ under a fixed number of steps of optimization will decrease with $\gamma_{0}$ at fixed learning rate. As a consequence, higher learning rates can be tolerated at larger widths. However, once the network is wide enough to be sufficiently close to the kernel regime, the maximum stable learning rate stabilizes as the dynamics become asymptotically linear. A comparison showing this trend is provided in 7.
@@ Appendix C Residual Networks
After these response functions have been solved for, we must solve for the correlation functions using the single site densities. This gives a closed form prediction for the loss dynamics.
@@ Appendix D Dynamics of Online SGD
$h^{0}(t)=u^{0}(t)+v(t)\ ,\ u^{0}(t)\sim\mathcal{N}\left(0,\frac{1}{\alpha_{B}}C_{\Delta}\ \delta(t-t^{\prime})\right)$ (87)
@@
The rest of the DMFT equations for $\{h^{\ell}(t),g^{\ell}(t)\}_{\ell\in\{1,..,L\}}$ are the same as before.

### v3

@@ 2 Deep Linear Network with Random Data
$f(\bm{x})=\frac{\sqrt{D}}{N\gamma_{0}}\left(\bm{w}^{L}\right)^{\top}\prod_{\ell=1}^{L-1}\left(\frac{1}{\sqrt{N}}\bm{W}^{\ell}\right)\left(\frac{1}{\sqrt{D}}\bm{W}^{0}\right)\bm{x}.$ (1)
The first matrix has dimensions $\bm{W}^{0}\in\mathbb{R}^{N\times D}$, while all intermediate weight matrices are all of size $\bm{W}^{\ell}\in\mathbb{R}^{N\times N}$. Lastly, the output $\bm{w}^{L}\in\mathbb{R}^{N}$ maps to a single output. We can provide extensions to unmatched hidden widths, structured data, and multiple (finite) output classes in Appendix F, E, and G, but we begin our analysis with this simplest possible case. We start by characterizing the dynamics for a random training dataset $\mathcal{D}=\{\bm{x}_{\mu},y_{\mu}\}_{\mu=1}^{P}$ of size $P$, and random initial weights
@@ 2.1 Dynamics of (S)GD in a Joint Asymptotic Limit
$\bm{v}(t)=\bm{w}^{\star}-\frac{\sqrt{D}}{N\gamma_{0}}\bm{W}^{0}(t)^{\top}\bm{g}^{1}(t).$ (9)
We stress that the number of these order parameters does not depend on $D,N,P$ but only scales with the number of iterations $T$ and the number of layers $L$, which justifies the saddle point derivation in the Appendix A. From the correlation functions, we can directly obtain the test loss $\mathcal{L}(t)$ and train loss ^$\hat{\mathcal{L}}(t)$ as
@@
Figure 2: The theory can capture the failure/success of learning rate transfer in NTK/$\mu$P networks. (a) The optimal learning rate increases with model width $\nu=N/D$ in NTK parameterization. This is due to the fact that the effective feature learning rate $\gamma_{0}$ is decreasing as $1/\sqrt{\nu}$. (b) The optimal learning rate is approximately preserved across $\nu=N/D$ in $\mu$P networks where $\gamma_{0}$ is constant, enabling learning rate transfer.
(a) Test Loss Varying Data $\gamma_{0}=2$
(b) Emergence of Overfitting from Data Repetition
(c) Online SGD Dynamics $\nu=0.25$
(d) SGD varying $\nu$ at Small $\alpha_{B}=0.05$
Figure 3: Train and test losses of a depth $L=4$ linear network versus dataset size $P=\alpha D$ compared to online SGD with batch size $B=\alpha_{B}D$. (a) Networks can fit the training data faster for smaller datasets $\alpha$. (b) The test error can be non-monotonic as a function of data $\alpha$ if training is performed too long. (c) The dynamics of online SGD with $\nu=0.5$ across batchsizes $B=\alpha_{B}D$. Smaller batchsizes slow down training through accumulated SGD variance, but there is not a bottleneck error as $t\to\infty$ provided the dynamics are stable. (d) Increasing the width $\nu$ eliminates the initial overfitting transient (see Appendix D.2).
@@ 3 Residual Networks and Depth Limits
$f(\bm{x})=\frac{\sqrt{D}}{N\gamma_{0}}\left(\bm{w}^{L}\right)^{\top}\prod_{\ell=1}^{L-1}\left(\bm{I}+\frac{\beta}{\sqrt{N}}\bm{W}^{\ell}\right)\left(\frac{1}{\sqrt{D}}\bm{W}^{0}\right)\bm{x}$ (16)
@@ 4 Structured Data & Power Law Scalings
which agrees with the exponents derived by Bordelon et al. (2024b) in the one-hidden layer $L=1$ case. However, our theory shows that this improved exponent for the hard task regime due to non-lazy dynamics is in fact the same for deeper networks as well. This indicates that the structure of the task could induce some universal behavior in linear networks of different architectures. We stress that this scaling law result is derived in the gradient flow limit of our more general DMFT, but edge of stability (EoS) effects could spoil this behavior at long enough training times in discrete time (Ren et al., 2025; Cohen et al., 2021). While this effect should in principle be availabel from our existing DMFT equations, we leave a more detailed analysis of this effect in this model to future work.
LR Transfer For Dimension Free Power Law Features
While our learning rate transfer plots exhibited very sharp dependence on learning rate for isotropic features (Figure 2), we can develop more realistic hyperparameter transfer curves under our power law feature setting. In this dimension-free setting ($D\to\infty$ with power law exponent $\alpha>1$), the theory depends explicitly on $N,B\gg 1$. We illustrate that our main findings regarding learning rate transfer persist for power law features in Figure 7.
(a) NTK Scaling
(b) $\mu$P Scaling
Figure 7: Learning rate transfer experiments for a $L=3$ $\mu$P linear network trained with SGD ($B=128$) on powerlaw features $(\alpha,\beta)=(2,1.75)$ and $\gamma_{0}=0.4$. (a) The hyperparameter transfer effect fails for NTK parameterization but (b) succeeds for $\mu$P and is accurately captured by DMFT (dashed).
@@ Discussion
The current theory, while a signficant improvement in complexity compared to prior works which do not average over the random data (such as Bordelon & Pehlevan (2022); Bordelon et al. (2024c)), becomes costly to compute for very deep networks trained for a large number of iterations $T$ and very deep networks $L\gg 1$. We summarize the memory and compute requirements to compute our theory in Table 1. Because of the $\sim L^{2}$ number of response functions required to compute our residual network DMFT, the costs are higher than for a standard non-residual network (which is linear in $L$). Our DMFT solver is more efficient than full batch training in a finite width $N$ linear ResNet provided that $LT^{2}\ll N^{2}P$. In addition, the function approximation power of this model is quite limited as we are only considering deep linear networks in the present study. However, despite the failure to capture nonlinear function approximation, many of the dynamical effects from initialization, SGD, and parameterization persist. Some future directions that could be analyzed within this framework include characterizing the impact of learning rate schedules and other optimizers on the dynamics, studying the effect of normalization layers, Hessian spectra, and characterizing how nonlinearity changes this theoretical picture.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgements
We thank Boris Hanin, Mufan (Bill) Li, Lorenzo Noci, Jacob Zavatone-Veth, Hugo Cui, Francesca Mignacco, Stefano Sarao Manelli, and Clarissa Lauditi for useful and engaging conversations. BB thanks Jason Lee for pointing out that discrete time effects in our power law model would alter the new scaling law behavior on hard tasks after sufficient training time.
@@ A.1.1 Isolating Dependence on The Disorder
$f(\bm{x})=\frac{\sqrt{D}}{N\gamma_{0}}(\bm{w}^{L})^{\top}\prod_{\ell=1}^{L-1}\left(\frac{1}{\sqrt{N}}\bm{W}^{\ell}\right)\left(\frac{1}{\sqrt{D}}\bm{W}^{0}\right)\bm{x}.$ (22)
@@
$\bm{w}^{L}(t+1)=\bm{w}^{L}(t)-N\eta\gamma_{0}^{2}\nabla_{\bm{w}^{L}}\hat{\mathcal{L}}(t)\ ,\ \bm{W}^{\ell}(t+1)=\bm{W}^{\ell}(t)-N\eta\gamma_{0}^{2}\nabla_{\bm{W}^{\ell}}\hat{\mathcal{L}}(t)\ ,\ \ell\in\{0,...,L-1\}.$ (24)
@@ A.2 Final DMFT Equations
$v(t)=w^{\star}-r^{0}(t)-\sum_{t^{\prime}<t}\left[\frac{1}{\gamma_{0}}R^{1}_{gu}(t,t^{\prime})+\eta\gamma_{0}C_{g}^{1}(t,t^{\prime})\right]h^{0}(t^{\prime})\ ,\ r^{0}(t)\sim\mathcal{GP}\left(0,\frac{1}{\nu\gamma_{0}^{2}}G^{1}\right)$
@@ A.3 Centering: Subtracting the Initial Predictor for Lazy Limit Stability
$v(t)=w_{\star}-r^{0}(t)+r^{0}(0)-\frac{1}{\gamma_{0}}\sum_{t^{\prime}<t}R^{1}_{gu}(t,t^{\prime})h^{0}(t^{\prime})-\eta\sum_{t^{\prime}<t}C_{g}^{1}(t,t^{\prime})h^{0}(t^{\prime})$
@@ Appendix B Comparing Parameterizations
In this section, we briefly comment on how the maximal stable learning rate scales with $\nu$ in NTK and in mean field/$\mu$P parameterization. NTK parameterization is defined as scaling the network so that $\gamma_{0}=\Theta\left(\frac{1}{\sqrt{\nu}}\right)$. However, a consequence of this choice is that the change in the correlation functions $C_{g}(t,t^{\prime})$ under a fixed number of steps of optimization will decrease with $\gamma_{0}$ at fixed learning rate. As a consequence, higher learning rates can be tolerated at larger widths. However, once the network is wide enough to be sufficiently close to the kernel regime, the maximum stable learning rate stabilizes as the dynamics become asymptotically linear. A comparison showing this trend is provided in 8.
@@ Appendix C Residual Networks
After these response functions have been solved for, we must solve for the correlation functions using the single site densities. Concretely, we use the above recursion to obtain the $\ell=k$ diagonal elements
Using the above response functions and invoking the linearity of the dynaimcs, we have the following decompositions of the $h^{\ell}(t),g^{\ell}(t)$ fields
$h^{\ell}(t)=\sum_{kt^{\prime}}\left[\frac{\partial h^{\ell}(t)}{\partial u^{k}(t^{\prime})}u^{k}(t^{\prime})+\frac{\partial h^{\ell}(t)}{\partial r^{k}(t^{\prime})}r^{k}(t^{\prime})\right]$ (76)
$g^{\ell}(t)=\sum_{kt^{\prime}}\left[\frac{\partial g^{\ell}(t)}{\partial u^{k}(t^{\prime})}u^{k}(t^{\prime})+\frac{\partial g^{\ell}(t)}{\partial r^{k}(t^{\prime})}r^{k}(t^{\prime})\right].$ (77)
We next compute the correlation and response functions $C^{\ell}_{h}(t,t^{\prime})$ and $C^{\ell}_{g}(t,t^{\prime})$ which have the form
$C^{\ell}_{h}(t,t^{\prime})=\sum_{kk^{\prime}ss^{\prime}}\frac{\partial h^{\ell}(t)}{\partial u^{k}(s)}\frac{\partial h^{\ell}(t^{\prime})}{\partial u^{k}(s^{\prime})}C^{k-1}_{h}(s,s^{\prime})+\sum_{kk^{\prime}ss^{\prime}}\frac{\partial h^{\ell}(t)}{\partial r^{k}(s)}\frac{\partial h^{\ell}(t^{\prime})}{\partial r^{k}(s^{\prime})}C^{k+1}_{g}(s,s^{\prime})$ (78)
$C^{\ell}_{g}(t,t^{\prime})=\sum_{kk^{\prime}ss^{\prime}}\frac{\partial g^{\ell}(t)}{\partial u^{k}(s)}\frac{\partial g^{\ell}(t^{\prime})}{\partial u^{k}(s^{\prime})}C^{k-1}_{h}(s,s^{\prime})+\sum_{kk^{\prime}ss^{\prime}}\frac{\partial g^{\ell}(t)}{\partial r^{k}(s)}\frac{\partial g^{\ell}(t^{\prime})}{\partial r^{k}(s^{\prime})}C^{k+1}_{g}(s,s^{\prime})$ (79)
where the base cases are determined by
$C^{0}_{h}(t,t^{\prime})=\sum_{ss^{\prime}}\mathcal{H}(t,s)\mathcal{H}(t^{\prime},s^{\prime})\left[1+\frac{1}{\nu\gamma_{0}^{2}}C^{1}_{g}(s,s^{\prime})\right]$ (80)
where the transfer function $\mathcal{H}(t,s)$ satisfies
$\mathcal{H}(t,s)=\delta(t-s)+\sum_{t^{\prime}<t}\left[\frac{1}{\gamma_{0}}R^{1}_{gu}(t,t^{\prime})+\eta C^{1}_{g}(t,t^{\prime})\right]\mathcal{H}(t^{\prime},s).$ (81)
The test loss can finally be obtained from the $t,t$ diagonal entries
$\mathcal{L}(t)=C_{h}^{0}(t,t).$ (82)
We see that the formulas for the response functions depend on the correlation functions and vice versa. These must all be solved for simultaneously.
@@ C.1 The Large $L$ Limit
Again from these response functions, we can again obtain the final correlation functions,
[…]
and the test loss dynamics have the form
$\mathcal{L}(t)=C_{h}(0,t,t).$ (95)
@@ Appendix D Dynamics of Online SGD
$h^{0}(t)=u^{0}(t)+v(t)\ ,\ u^{0}(t)\sim\mathcal{N}\left(0,\frac{1}{\alpha_{B}}C_{\Delta}(t,t^{\prime})\ \delta(t-t^{\prime})\right)$ (97)
@@
The rest of the DMFT equations for $\{h^{\ell}(t),g^{\ell}(t)\}_{\ell\in\{1,..,L\}}$ are the same as before as we provide below
D.1 Full DMFT Equations for Online SGD
Below we provide the full set of DMFT equations for the online learning setting where batchsize $B$, width $N$ and dimension $D$ diverge proportionally
$D,N,B\to\infty\ ,\ B=\alpha D\ ,\ N=\nu D.$ (99)
In this limit, we obtain the following stochastic process for $\{v(t),\Delta(t),\{h^{\ell}(t)\}_{\ell=0}^{L},\{g^{\ell}(t)\}_{L=1}^{L}\}$
$v(t)=w^{\star}-r^{0}(t)-\sum_{t^{\prime}<t}\left[\frac{1}{\gamma_{0}}R^{1}_{gu}(t,t^{\prime})+\eta\gamma_{0}C_{g}^{1}(t,t^{\prime})\right]h^{0}(t^{\prime})\ ,\ r^{0}(t)\sim\mathcal{GP}\left(0,\frac{1}{\nu\gamma_{0}^{2}}G^{1}(t,t^{\prime})\right)$
$\Delta(t)=u_{\Delta}(t)+\sigma\epsilon(t)\ ,\ u_{\Delta}(t)\sim\mathcal{GP}\left(0,C_{v}(t.t^{\prime})\right)\ ,\ \epsilon(t)\sim\mathcal{N}(0,\delta(t-t^{\prime}))$
$h^{0}(t)=u^{0}(t)+v(t)\ ,\ u^{0}(t)\sim\mathcal{GP}\left(0,\frac{1}{\alpha_{B}}C_{\Delta}(t,t^{\prime})\delta(t-t^{\prime})\right)$
$h^{1}(t)=u^{1}(t)+\sum_{t^{\prime}<t}\left[\frac{1}{\nu\gamma_{0}}R_{hr}^{0}(t,t^{\prime})+\eta\gamma_{0}C_{h}^{0}(t,t^{\prime})\right]g^{1}(t^{\prime})\ ,\ u^{1}(t)\sim\mathcal{GP}(0,C_{h}^{0}(t,t^{\prime}))$
$h^{\ell}(t)=u^{\ell}(t)+\sum_{t^{\prime}<t}\left[R_{hr}^{\ell-1}(t,t^{\prime})+\eta\gamma_{0}C_{h}^{\ell-1}(t,t^{\prime})\right]g^{\ell}(t^{\prime})\ ,\ u^{\ell}(t)\sim\mathcal{GP}(0,C_{h}^{\ell-1}(t,t^{\prime}))$
$g^{\ell}(t)=r^{\ell}(t)+\sum_{t^{\prime}<t}\left[R_{gu}^{\ell+1}(t,t^{\prime})+\eta\gamma_{0}C_{g}^{\ell+1}(t,t^{\prime})\right]h^{\ell}(t^{\prime})\ ,\ r^{\ell}(t)\sim\mathcal{GP}(0,C_{g}^{\ell+1}(t,t^{\prime}))$ (100)
where as before the response functions are defined as
$R_{vu}^{0}(t,t^{\prime})=\left<\frac{\partial v(t)}{\partial u^{0}(t^{\prime})}\right>\ ,\ R^{0}_{hr}(t,t^{\prime})=\left<\frac{\partial h^{0}(t)}{\partial r^{0}(t^{\prime})}\right>$
$R^{\ell}_{hr}(t,t^{\prime})=\left<\frac{\partial h^{\ell}(t)}{\partial r^{\ell}(t^{\prime})}\right>\ ,\ R^{\ell}_{gu}(t,t^{\prime})=\left<\frac{\partial g^{\ell}(t)}{\partial u^{\ell}(t^{\prime})}\right>,$ (101)
and the correlation functions are defined as usual
$C_{h}^{\ell}(t,t^{\prime})=\left<h^{\ell}(t)h^{\ell}(t^{\prime})\right>\ ,\ C_{g}^{\ell}(t,t^{\prime})=\left<g^{\ell}(t)g^{\ell}(t^{\prime})\right>.$ (102)
D.2 Analysis of Early Portion of Online SGD
From our theory and simulations, we see that for sufficiently large learning rates and sufficiently small width and batch size, the test loss can initially increase with iterations $t$ before subsequently recovering and converging at later time. In this section, we analyze this effect as an approximation of the early loss dynamics. Neglecting the memory terms that buildup over longer training times, we make the following approximations for small $t$
$R_{gu}^{1}(t,t^{\prime})\approx\eta\gamma_{0}L\Theta(t-t^{\prime})$
$G^{1}(t+1,t+1)-G^{1}(t,t)\approx[G^{2}(t+1,t+1)-G^{2}(t,t)]+4\eta^{2}L^{2}\gamma_{0}^{2}C_{h}^{0}(t,t)$
$\approx\eta^{2}C_{h}^{0}(t,t)\sum_{k=1}^{L}k^{2}=\frac{\eta^{2}\gamma_{0}^{2}L(L+1)(2L+1)}{6}C_{h}^{0}(t,t)$ (103)
which can be justified by working out the early dynamics for the response functions at small time. Under this approximation, we derive the following recursion for the test loss $C_{v}(t)=\left<v(t)^{2}\right>$, which has the form
$C_{v}(t+1)\approx\left[\left(1-L\eta\right)^{2}+\frac{\eta^{2}L^{2}}{\alpha_{B}}+\frac{\eta^{2}L(L+1)(2L+1)}{6\nu}+\frac{\eta^{2}L(L+1)(2L+1)}{6\nu\alpha_{B}}\right]C_{v}(t).$ (104)
We verify that this does accurately predict the first few steps of the loss dynamics (and captures the early blowup) of deep linear networks in Figure 9. Under this early time approximation, the loss will thus increase if
$\eta L>\frac{2}{1+\frac{1}{\alpha_{B}}+\frac{(L+1)(2L+1)}{6L\nu}+\frac{(L+1)(2L+1)}{6L\alpha_{B}\nu}}.$ (105)
Indeed from simulations, we see that the early dynamics diverge for sufficiently small $\alpha_{B}$ and $\nu$.
(a) $L=4,\nu=0.25$
(b) $L=4,\alpha_{B}=0.05$
Figure 9: Early dynamics of the online SGD can exhibit an initial blowup followed by recovery if $\gamma_{0}>0$. The initial blowup is approximately $\mathcal{L}(t)\approx\left[(1-\eta L)^{2}+\eta^{2}L^{2}\left(\alpha_{B}^{-1}+\nu^{-1}+2\alpha_{B}^{-1}\nu^{-1}\right)\right]^{t}$, which is a function of learning rate $\eta$, depth $L$, batchsize $\alpha_{B}$ and width $\nu$. Small width or small batchsize can make this effect more pronounced.